If the sum of p,q,r terms of an a.p are a,b,c respectively then prove: a/p (q-r)+b/p (r-p) +c/r (p-q)=0 . Anybody plzz do this ques

Grade:10

1 Answers

Arun
6 years ago
Dear Vaibhav
 
let first term = x   and common diff. = y
 
now
a = (p/2) [2x + (p – 1)y]
hence
a/p = x + (p –1)y/2
 
similarly
b/q = x + ( q–1)y/2
&
c/r = x + (r –1)y/2
 
hence put these in the LHS of equation.
 
you will simply get your result as zero
 
Regards
Arun (askIITians forum expert)